OpenStudy (anonymous):

I am not sure how to ask this 1. I need to find an equation of the line L. y=-x and L is perpendicular. y |3 | (-1,1)* | -|-|-|-|-|-|x -3 | 3

OpenStudy (anonymous):

y=-x has a slope of -1/1 A perpendicular line has a slope that is a negative reciprocal, so slope=1/1=1 If the line goes through point (-1,1) then we can substitute its x and y values, as well as our new found slope into our general equation of y=mx+b to find its y-int \[(1)=(1)(-1)+b\] \[(1)=(-1)+b\] \[2=b\] so our equation is: \[y=(1)x+2\] or \[y=x+2\]

OpenStudy (anonymous):

Thank you so much for your help.

OpenStudy (anonymous):

Awesome job Mr. Yantho! Great question kcbrosell

OpenStudy (anonymous):

Thank you